OpenStudy (anonymous):

A carpenter is putting a skylight in a roof. If the roof measures 10x + 9 by 7x + 7 and the skylight measures x + 5 by 3x + 3, what is the area of the remaining roof after the skylight is built?

OpenStudy (campbell_st):

area of the roof is A = (10x +9)(7x + 7) area of the skylight is S=(x + 5)(3x + 3) to the area remaining =(10x +9)(7x +7) - (x +5)(3x + 3)

OpenStudy (anonymous):